OpenStudy (anonymous):

What is the range of this? h(x) = 2x^2+x -1?

OpenStudy (anonymous):

\[[-1,+ \infty)\]

OpenStudy (amistre64):

range is, by convention, determined by the highest and lowest points; and excluding what nneds to be excluded

OpenStudy (amistre64):

since this is an equation of a parabola that opens upward; we only need to determine the vertex position which will be the lowest point

OpenStudy (amistre64):

the vertex can be gotten from the determinant of the quadratic formula: -b/2a will be the x part

OpenStudy (amistre64):

ax^2 +bx+c is the general quadratic; from this we see that -1/4 is the determinant of your equation

OpenStudy (amistre64):

2(-1/4)^2 -(1/4)-1 = -9/8 for the lowest y value in the range

OpenStudy (amistre64):

therefore our range is [-9/8,inf)

OpenStudy (angela210793):

y of the vertex is =-D/4a isn't it?

OpenStudy (amistre64):

y is determined by plugging "D" back into the equation and solving

OpenStudy (amistre64):

-1/4 ----- not = -9/8 so id venture to say no :) 4 (2)

OpenStudy (angela210793):

D=b^2-4ac=9 tht's wht i am talking abt

OpenStudy (anonymous):

r-(-9/8)